OpenStudy (anonymous):

suppose you flip a coin n times, and the probability of getting heads 15 times is .0148. What's n? A. 19 B. 20 C. 16 D. 17 E. 24 I don't know how you would work backwards to find n help

OpenStudy (dumbcow):

use binomial theorem \[\left(\begin{matrix}n \\ 15\end{matrix}\right) (\frac{1}{2})^{15} (\frac{1}{2})^{n-15} = 0.0148\] which reduces to \[\left(\begin{matrix}n \\ 15\end{matrix}\right) (\frac{1}{2})^n = .0148\] this cannot be solved algebraically, but since its multiple choice just plug in each option for "n"

OpenStudy (anonymous):

oh okay so you would divide n by 15 right and then multipy by the 1/2^n? @dumbcow

OpenStudy (anonymous):

im also getting all these numbers like 1.2768352e-6? for everyone am i doing something wrong?

OpenStudy (dumbcow):

no its combinations not division' \[\left(\begin{matrix}n \\ 15\end{matrix}\right) = \frac{n!}{(n-15)! 15!}\]

OpenStudy (anonymous):

okay but for some reason im still getting everything to the e power so i think im doing something wrong.